A quantum particle confined to one dimensional box of width'a' is in its first excited state. What is the probability of finding the particle over an interval of (a/2) marked symmetrically at the centre of the box
1
Expert's answer
2021-08-16T04:56:01-0400

"\\psi=\\sqrt\\frac{2}a sin (\\frac{\u03c0nx}{a})\n\u200b\t\n \n\u200b"

"n=2"